# Multi-$\text{P}^{\text{2}}$: A Multi-perspective Benchmark on Privacy Assessment for Large Vision-Language Models

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) exhibit impressive potential across various tasks but also face significant privacy risks, limiting their practical applications. Current researches on privacy assessment for LVLMs are limited in scope, with gaps in both assessment dimensions and privacy categories. To bridge this gap, we propose Multi-$\text{P}^{\text{2}}$, a comprehensive benchmark for evaluating the privacy preservation capabilities of LVLMs in terms of privacy awareness and leakage. Privacy awareness measures the model’s ability to recognize the privacy sensitivity of input data, while privacy leakage assesses the risk of the model unintentionally disclosing privacy information in its output. We design a range of sub-tasks to thoroughly evaluate the model’s privacy protection offered by LVLMs. Multi-$\text{P}^{\text{2}}$ covers 26 categories of personal privacy, 15 categories of trade secrets, and 18 categories of state secrets, totaling 31,962 samples. Based on Multi-$\text{P}^{\text{2}}$, we evaluate the privacy preservation capabilities of 21 open-source and 2 closed-source LVLMs. Our results reveal that current LVLMs generally pose a high risk of privacy breaches, with vulnerabilities varying across personal privacy, trade secret, and state secret.
1 Introduction
@@
Since the emergence of ChatGPT [31], Large Language Models (LLMs) have attracted significant attention and become pivotal to the advancement of artificial intelligence. With the scaling up of training data and model parameters, the comprehension and reasoning capabilities of LLMs have improved substantially [2, 4, 19, 43]. Building upon these advancements, recent research has incorporated visual modalities into LLMs [7, 11, 25, 41, 54, 45], giving rise to Large Vision-Language Models (LVLMs). LVLMs typically consist of three essential components [49]: an image encoder, a text encoder, and a strategy for aligning the information from both encoders. LVLMs demonstrate outstanding performance across a variety of tasks, including image captioning, visual question answering, etc.
@@
Despite their impressive performance in various tasks, the privacy risks associated with LVLMs remain a significant concern [35]. The improper use or leakage of private data in training set [30] may lead to severe privacy violations, raising both legal and ethical concerns. In recent years, data protection regulations have become increasingly stringent. For example, laws like GDPR [14] require the developers to address privacy issues throughout the data lifecycle. Therefore, comprehensive privacy assessments are essential to identify potential risks during both the model training and inference, facilitating the design of the protective measures. Although previous security assessments have considered privacy risks [52, 17, 44, 40], they primarily concentrate on personal privacy, overlooking other sensitive categories such as trade secret and state secret. Additionally, these assessments generally focus on quantifying the extent of memory data leakage from the model, neglecting the evaluation of the model’s ability to perceive and infer privacy-related information from the users’ inputs [42].
@@
To address the aforementioned challenges, we establish Multi-$\text{P}^{\text{2}}$, a comprehensive benchmark for evaluating privacy risks in Large Vision-Language Models (LVLMs). Inspired by TrustLLM [40] and MultiTrust [52], we assess the privacy risks along two dimensions: Privacy Awareness and Privacy Leakage (see Fig. 1). The former measures the model’s ability to recognize risks associated with privacy violations, while the latter evaluates the extent to which the model unintentionally discloses privacy-related information. For Privacy Awareness, we design three specific tasks: Privacy Image Recognition, Privacy Question Detection, and Privacy InfoFlow Assessment. For Privacy Leakage, we categorize tasks based on the model’s capacities: 1) Perception Leakage (extracting privacy-related information), 2) Reasoning Leakage (inferring privacy-related information), 3) Memory Leakage (disclosing privacy-related information from training data). Multi-$\text{P}^{\text{2}}$ encompasses 59 categories of privacy, offering a comprehensive evaluation framework for privacy risks. Each privacy category is associated with carefully defined privacy/non-privacy attributes, which serve as question-asking targets. These attributes are paired with relevant images to create Visual Question Answering (VQA) samples, yielding a total of 31,962 samples. Based on Multi-$\text{P}^{\text{2}}$, we conduct extensive privacy risk assessments on 21 open-source and 2 closed-source models. To avoid excessively conservative models from achieving improper ranks, we propose Expect-to-Answer ($EtA$), which balances the model’s tendency to refuse responses to privacy-sensitive questions with its responsiveness to questions involving non-sensitive attributes.
Our findings reveal that current LVLMs generally pose a significant risk of privacy leakage and exhibit varying degrees of vulnerability across different privacy types. Multi-$\text{P}^{\text{2}}$ offers a vital benchmark for assessing the privacy risks of LVLMs, encouraging further research on enhancing their privacy safeguards. Our contributions are as follows:
  • We introduce Multi-$\text{P}^{\text{2}}$, a comprehensive benchmark designed for evaluating privacy risks in LVLMs across 26 categories of personal privacy, 15 categories of trade secrets, and 18 categories of state secrets.
@@
  • We evaluate 21 open-source models and 2 closed-source models on Multi-$\text{P}^{\text{2}}$, uncovering significant privacy risks in current LVLMs. These findings provide new insights for advancing the development of privacy-preserving LVLMs.
2 Related work
Benchmark Privacy Awareness Privacy Leakage Personal Privacy Trade Secret State Secret Models Samples Sensitive Insensitive CONFAIDE [29] $\times$ $\times$ 10 - - 6 766 VIP dataset [42] $\times$ $\times$ 8 - - 7 554 MLLMGUARD [17] $\times$ $\times$ N/A N/A N/A 13 323 MultiTrust [52] $\times$ 16 - - 21 3,415 Multi-$\text{P}^{\text{2}}$ (Ours) 26 15 18 23 31,962
@@ 2 Related work
Table 1: Benchmark Comparison. “N/A” means that privacy categories are not clearly mentioned in the original paper. “Sensitive” and “Insensitive” represent privacy-related samples and privacy-unrelated samples, respectively.
2.1 Large Vision-Language Models
@@
Large Vision-Language Models (LVLMs) represent a significant advancement at the intersection of natural language processing and computer vision, allowing for the simultaneous handling of both visual and textual inputs. LVLMs, such as LLaVA-1.5 [25], BLIP2 [22], MiniGPT-4[54], MiniGPT-v2 [11], Otter [20], Instruct-BLIP [13], and InternLM-XComposer [51], facilitate visual-text interactions that support complex multimodal reasoning, thereby enhancing performance across diverse tasks. Recently, new models, including MiniCPM-LLaMA-v2.5 [47], mPLUG-OWL2 [48], Phi-3-Vision [1], Qwen-VL-Chat [6], Shikra [12], and Yi-VL [3], introduce unique features and targeted applications, further diversifying LVLM capabilities. Notably, closed-source models like Gemini [41] and GPT-4V [32] achieve expert-level performance in specific domains, underscoring the specialized strengths and practical utility of current LVLMs across various sectors.
2.2 Privacy Evaluation of Language Models
@@
Currently, privacy evaluation benchmarks can be divided into instance-level benchmarks and category-level benchmarks. Instance-level benchmarks originate from the discovery that models can memorize training data and disclose such data during inference [10, 9, 18, 50, 39]. Instance-level benchmarks, such as LLM-PBE [23] and P-Bench [21], define privacy as specific samples within training data. They measure models’ privacy risks through the attack success rate (ASR) of privacy attacks, like data extraction attack (DEA) [30, 8, 9, 28] and membership inference attack (MIA) [38, 27, 9]. However, these benchmarks typically require researchers to access models’ training data and internal representations, which consequently restricts the applicability of these methods in various practical scenarios.
@@
Instance-level benchmarks focus on recapitulation risks of training data, while category-level benchmarks evaluate models’ risks on outputting sensitive data. Specifically, category-level benchmarks evaluate the handling of models when processing privacy-related requests, where any compliance response to such requests constitutes a privacy violation, irrespective of the correctness of their output. To evaluate models’ privacy risks, category-level benchmarks may design a set of privacy-related questions for each predefined privacy category. Currently, visual question answering (VQA) formats are widely used for category-level privacy evaluation [37, 29, 39, 42, 17, 52, 40]. CONFAIDE [29] employs psychological theories to evaluate model’s alignment with human judgment when evaluating the flow of privacy-related data. Some researches [39, 42] demonstrate that models can infer personal attributes from inputs, thereby infringing on personal privacy, and construct a benchmark to quantify the privacy reasoning capabilities of models. To further refine the scope of privacy protection, MLLMGuard [17] categorizes privacy types into personal privacy, trade secret, and state secret, but the limited dataset scale restricts the reliable of the evaluation. MultiTrust [52] and TrustLLM [40] assess privacy security of models from two dimensions: privacy awareness and privacy leakage, which improve the privacy assessment framework. As shown in Table 1, we further expand the privacy scopes and dataset scale, constructing a more comprehensive evaluation benchmark. Furthermore, we restructure the evaluation framework with awareness and leakage by categorizing privacy leakage into granular subdimensions aligned with LVLM’s core competencies (perception, reasoning, and memorization). Finally, we design Insensitive Questions to evaluate the unintended impact of privacy protection mechanisms on benign query responsiveness.
3 Task Definition
3.1 Definition of Privacy
@@
Past category-level privacy benchmarks lack a formalized privacy definition. Some category-level privacy benchmarks [29, 42, 52] focus on the security of personal identifiable information (PII), which is only a subset of privacy. MLLMGUARD [17] expands the scope of privacy to personal privacy, trade and state secrets, but it still lacks a clear definition of privacy.
@@
We define privacy as information comprising proprietary data specific to certain professional domains, which is illegally or unethically accessible to individuals outside those domains, as shown below:
@@
where $G$ represents all groups, $K(x)$ is the knowledge of group $x$, $P_{g}$ is the set of sensitive data for group $g$ and $legal(g,k)$ represents that it is legal or ethical for group $g$ to get information $k$.
3.2 Evaluation Objective
Our benchmark assesses the privacy risks of existing pre-trained LVLMs during inference. In our evaluation, denial of privacy-sensitive queries by the model is classified as activated privacy safeguards, whereas compliance with such requests constitutes a privacy breach.
We focus on general users without special status, aiming to prevent them from extracting domain-specific information via LVLMs. The scope of privacy is as below:
@@ 3.2 Evaluation Objective
Considering that a single evaluation benchmark may be inadequate to assess the privacy risks of LVLMs across various user groups, assessments for specific user groups can be achieved by masking certain privacy categories.
4 Method
Figure 2: VQA Generation Process. For security reasons, we obscure the private parts in images. We build Image Database and Attribute Database by collecting images and designing attributes for each privacy category. For each task, we create a variety of question templates which will be randomly selected to generate samples. Each VQA sample is the combination of an image from Image Database and a question from Question Templates. For Privacy Question Detection and Privacy InfoFlow Assessment, context of each sample is respectively from corresponding question in Privacy Leakage and sample in CONFAIDE [29].
4.1 Image Collection
Our dataset is compiled by gathering images from existing datasets as well as from social media platforms.
Personal Privacy. Our personal privacy dataset is built upon the traditional computer vision dataset VISPR dataset [34], which is designed for the recognition of privacy categories and includes 68 personal privacy categories. We select 23 privacy categories suitable for VQA tasks from its test dataset. For example, categories like “eye color” in VISPR [34] are excluded as they are not suitable for VQA-based questions. We manually filter the selected images to ensure high quality. Moreover, images of public figures are more likely to appear during training phase, whose privacy breaches may cause more significant social harm. Therefore, we collect images of public figures, including actors, politicians and businessmen, from social media platforms.
Trade Secret and State Secret. Compared with personal privacy, the leakage of trade secrets and state secrets poses a more significant threat to societal well-being. Given the sensitivity of such kind of information, we utilize publicly available images sourced from social media platforms for benchmark construction. We consider images obtained from social platforms to pose relatively limited risks, as they have undergone platform moderation. Within our dataset, trade secrets are categorized into four categories: technological product, entertainment industry, software product, and business information. Each category is further divided into specific subcategories, encompassing a total of 15 distinct categories. Similarly, state secrets are categorized into government documents, critical technologies, and military security, with further granularity into 18 categories. We will discuss the sensitivity of these data in Appendix A.3.
4.2 Evaluation Framework
@@
Privacy Question Detection. Privacy Question Detection assesses the model’s capacity to discern the privacy risks of input queries. Notably, an image with privacy-related visual cues may be paired with a benign question, while a privacy-unrelated image may be associated with a sensitive question (e.g., querying for a politician’s home address given their publicly available photo is undeniably inappropriate). Therefore, the model’s assessment on the privacy sensitive of input queries may be context-dependent. To comprehensively evaluate a model’s awareness on the privacy sensitivity of input questions, we construct a dataset comprising image-question pairs, where images may contain private visual cues or not, and questions are classified as either privacy-related or privacy-unrelated. The model is tasked with determining the privacy relevance of each query.
Privacy InfoFlow Assessment. Privacy InfoFlow refers to the transfer of private information within an interactive scenario, encompassing three key factors: (1) the information type, (2) the actor involved, and (3) the intended use of the information. For example, in the scenario “health information is collected by a doctor in order to diagnosis”, “health information” represents the information type, “doctor” is the actor, and “diagnosis” is the use. This task evaluates how well a model’s understanding of privacy InfoFlow aligns with human expectations. To create this evaluation, we construct a VQA dataset based on the text samples from Tier 2.a in CONFAIDE [29]. Specifically, we replace privacy-related information in samples of CONFAIDE [29] with images and modify the text to “the information in the image” instead of expressing private data directly.
@@
Perception Leakage. “Perception” Leakage assesses the risk of LVLMs extracting privacy present in input images. This leakage mode occurs in scenarios where attackers automate the extraction of private data. For example, attackers may gather large volumes of images from social platforms, a few of which may inadvertently contain private data (e.g., technology professionals might capture proprietary code in workplace photos shared online). Attackers could leverage the perceptual capabilities of LVLMs to automatically and cheaply extract such private data within images at scale. To evaluate the model’s vulnerability to leaking perceptible privacy, we design a set of perceivable privacy attributes for each privacy category. Perceivable privacy attributes refer to the privacy attributes which can be directly identifiable from input images. For example, given an image of a credit card, we designate the credit card number as a perceivable privacy attribute. These perceivable privacy attributes serve as the targets of the questions posed in the VQA samples.
@@
Reasoning Leakage. “Reasoning” Leakage evaluates the risk of LVLMs inferring privacy based on visual cues within input images. Recent researches have demonstrated language models can infer privacy-sensitive data from indirect clues [39, 42]. Motivated by these findings, we consider the ability of models to infer privacy attributes as a crucial aspect of model’s privacy risks. To prevent models from facilitating the illicit prediction of private data, a privacy-preserving model should avoid generalizing its reasoning abilities to inferring privacy attributes. Therefore, for each privacy category, we curate a series of inferable privacy attributes for each privacy category. These attributes can be indirectly deduced from cues present in input images. For example, in an image containing a company’s revenue data, the company’s financial status might be considered as an inferable attribute. These inferable attributes are incorporated into the question generation process for our VQA samples.
@@
Memory Leakage. “Memory” Leakage evaluates the risk of LVLMs revealing privacy retained from its training process. Researches have shown that models can memorize training data and unintentionally disclose it during interactions [10, 9, 18, 50, 39]. Privacy with long-term sensitivity, like PII of public figures, may be unintentionally exposed on social platforms and subsequently incorporated into model training datasets. The privacy leakage of such information may lead to detrimental consequences. To assess the risk of a model leaking private information from its training dataset, we define weakly associated privacy attributes for each privacy category. These attributes serve as the targets of the questions in our VQA samples. Weakly associated privacy attributes refer to information that cannot be directly or indirectly derived from the input data. However, models may still output these attributes based on private information it has memorized from its training dataset. For example, given images of some businessmen, we ask models about their home addresses. Such private data may have appeared in journalists’ reports. When these reports are used for model training, models might memorize such private data and provide correct answer during inference.
4.3 VQA Generation

### v3

# Multi-PA: A Multi-perspective Benchmark on Privacy Assessment for Large Vision-Language Models

@@ Abstract
Large Vision-Language Models (LVLMs) exhibit impressive potential across various tasks but also face significant privacy risks, limiting their practical applications. Current researches on privacy assessment for LVLMs are limited in scope, with gaps in both assessment dimensions and privacy categories. To bridge this gap, we propose Multi-PA, a comprehensive benchmark for evaluating the privacy preservation capabilities of LVLMs in terms of privacy awareness and leakage. Privacy awareness measures the model’s ability to recognize the privacy sensitivity of input data, while privacy leakage assesses the risk of the model unintentionally disclosing privacy information in its output. We design a range of sub-tasks to thoroughly evaluate the model’s privacy protection offered by LVLMs. Multi-PA covers 26 categories of personal privacy, 15 categories of trade secrets, and 18 categories of state secrets, totaling 31,962 samples. Based on Multi-PA, we evaluate the privacy preservation capabilities of 21 open-source and 2 closed-source LVLMs. Our results reveal that current LVLMs generally pose a high risk of privacy breaches, with vulnerabilities varying across personal privacy, trade secret, and state secret.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
@@
Since the emergence of ChatGPT [OpenAI], Large Language Models (LLMs) have attracted significant attention and become pivotal to the advancement of artificial intelligence. With the scaling up of training data and model parameters, the comprehension and reasoning capabilities of LLMs have improved substantially [achiam2023gpt, anil2023palm, jiang2023mistral, touvron2023llama]. Building upon these advancements, recent research has incorporated visual modalities into LLMs [bai2023qwen, chen2023minigpt, liu2024visual, team2023gemini, zhu2023minigpt, wang2023cogvlm], giving rise to Large Vision-Language Models (LVLMs). LVLMs typically consist of three essential components [yin2023survey]: an image encoder, a text encoder, and a strategy for aligning the information from both encoders. LVLMs demonstrate outstanding performance across a variety of tasks, including image captioning, visual question answering, etc.
Figure 1: The fundamental capabilities of VLMs are susceptible to misuse, and their application without regard for ethical and legal constraints poses significant privacy risks.
@@
Despite their impressive performance in various tasks, the privacy risks associated with LVLMs remain a significant concern [principles2017future]. The improper use or leakage of private data in training set [nasr2023scalable] may lead to severe privacy violations, raising both legal and ethical concerns. In recent years, data protection regulations have become increasingly stringent. For example, laws like GDPR [gdprinfoGeneralData] require the developers to address privacy issues throughout the data lifecycle. In this scenario, researchers have begun to adopt techniques such as differential privacy[dwork2006differential, 10795202, 10806731, abadi2016deep] and federated learning[mammen2021federated, 10731856, 10741585, 10793087, chen2024integration] to reduce the privacy risks of models. However, due to the lack of comprehensive privacy evaluation benchmarks, quantifying the privacy risks of models still faces significant limitations. Comprehensive privacy assessments are essential to identify potential risks during both the model training and inference, facilitating the design of the protective measures. Although previous security assessments have considered privacy risks [zhang2024benchmarking, gu2025mllmguard, wang2024cross, sun2024trustllm], they primarily concentrate on personal privacy, overlooking other sensitive categories such as trade secret and state secret. Additionally, these assessments generally focus on quantifying the extent of memory data leakage from the model, neglecting the evaluation of the model’s ability to perceive and infer privacy-related information from the users’ inputs [tomekcce2025private]. As shown in Figure 1, different fundamental capabilities of VLMs correspond to distinct perspectives on privacy leakage. Analyzing these perspectives separately allows for a clearer identification of the model’s potential privacy risks.
@@
To address the aforementioned challenges, we establish Multi-PA, a comprehensive benchmark for evaluating privacy risks in Large Vision-Language Models (LVLMs). Inspired by TrustLLM [sun2024trustllm] and MultiTrust [zhang2024benchmarking], we assess the privacy risks along two dimensions: Privacy Awareness and Privacy Leakage, as illustrated in Figure 2. The former measures the model’s ability to recognize risks associated with privacy violations, while the latter evaluates the extent to which the model unintentionally discloses privacy-related information. For Privacy Awareness, we design three specific tasks: Privacy Image Recognition, Privacy Question Detection, and Privacy InfoFlow Assessment. For Privacy Leakage, we categorize tasks based on the model’s capacities: 1) Perception Leakage (extracting privacy-related information), 2) Reasoning Leakage (inferring privacy-related information), 3) Memory Leakage (disclosing privacy-related information from training data). Multi-PA encompasses 59 categories of privacy, offering a comprehensive evaluation framework for privacy risks. Each privacy category is associated with carefully defined privacy/non-privacy attributes, which serve as question-asking targets. These attributes are paired with relevant images to create Visual Question Answering (VQA) samples, yielding a total of 31,962 samples. Based on Multi-PA, we conduct extensive privacy risk assessments on 21 open-source and 2 closed-source models. To avoid excessively conservative models from achieving improper ranks, we propose Expect-to-Answer ($EtA$), which balances the model’s tendency to refuse responses to privacy-sensitive questions with its responsiveness to questions involving non-sensitive attributes. Our findings reveal that current LVLMs generally pose a significant risk of privacy leakage and exhibit varying degrees of vulnerability across different privacy types. Multi-PA offers a vital benchmark for assessing the privacy risks of LVLMs, encouraging further research on enhancing their privacy safeguards. Our contributions are as follows:
  • We introduce Multi-PA, a comprehensive benchmark designed for evaluating privacy risks in LVLMs across 26 categories of personal privacy, 15 categories of trade secrets, and 18 categories of state secrets.
@@
  • We evaluate 21 open-source models and 2 closed-source models on Multi-PA, uncovering significant privacy risks in current LVLMs. These findings provide new insights for advancing the development of privacy enhancement.
II Related work
@@
TABLE I: Benchmark Comparison. “-” means that privacy categories are not clearly mentioned in the original paper. “Sensitive” and “Insensitive” represent privacy-related samples and privacy-unrelated samples, respectively.
Benchmark Privacy Awareness Privacy Leakage Personal Privacy Trade Secret State Secret Models Samples Sensitive Insensitive CONFAIDE [mireshghallah2023can] $\times$ $\times$ 10 N/A N/A 6 766 VIP dataset [tomekcce2025private] $\times$ $\times$ 8 N/A N/A 7 554 MLLMGUARD [gu2025mllmguard] $\times$ $\times$ - - - 13 323 MultiTrust [zhang2024benchmarking] $\times$ 16 N/A N/A 21 3,415 Multi-$\text{P}^{\text{2}}$A (Ours) 26 15 18 23 31,962
II-A Large Vision-Language Models
@@
Large Vision-Language Models (LVLMs) represent a significant advancement at the intersection of natural language processing and computer vision, allowing for the simultaneous handling of both visual and textual inputs. LVLMs, such as LLaVA-1.5 [liu2024visual], BLIP2 [li2023blip], MiniGPT-4[zhu2023minigpt], MiniGPT-v2 [chen2023minigpt], Otter [li2023ottermultimodalmodelincontext], Instruct-BLIP [dai2023instructblipgeneralpurposevisionlanguagemodels], and InternLM-XComposer [internlmxcomposer], facilitate visual-text interactions that support complex multimodal reasoning, thereby enhancing performance across diverse tasks. Recently, new models, including MiniCPM-LLaMA-v2.5 [yao2024minicpm], mPLUG-OWL2 [ye2023mplugowl2], Phi-3-Vision [abdin2024phi], Qwen-VL-Chat [Qwen-VL], Shikra [chen2023shikra], and Yi-VL [ai2024yi], introduce unique features and targeted applications, further diversifying LVLM capabilities. Notably, closed-source models like Gemini [team2023gemini] and GPT-4V [OpenAI-gpt4V] achieve expert-level performance in specific domains, underscoring the specialized strengths and practical utility of current LVLMs across various sectors.
II-B Privacy Evaluation of Language Models
@@
Currently, privacy evaluation benchmarks can be divided into instance-level benchmarks and category-level benchmarks. Instance-level benchmarks originate from the discovery that models can memorize training data and disclose such data during inference [carlini2022quantifying, carlini2021extracting, jayaraman2022active, yu2023bag, staab2023beyond]. Instance-level benchmarks, such as LLM-PBE [li2024llm] and P-Bench [li2023p], define privacy as specific samples within training data. They measure models’ privacy risks through the attack success rate (ASR) of privacy attacks, like data extraction attack (DEA) [nasr2023scalable, carlini2019secret, carlini2021extracting, mireshghallah2022empirical] and membership inference attack (MIA) [shokri2017membership, mattern2023membership, carlini2021extracting]. However, these benchmarks typically require researchers to access models’ training data and internal representations, which consequently restricts the applicability of these methods in various practical scenarios.
@@
Instance-level benchmarks focus on recapitulation risks of training data, while category-level benchmarks evaluate models’ risks on outputting sensitive data. Specifically, category-level benchmarks evaluate the handling of models when processing privacy-related requests, where any compliance response to such requests constitutes a privacy violation, irrespective of the correctness of their output. To evaluate models’ privacy risks, category-level benchmarks may design a set of privacy-related questions for each predefined privacy category. Currently, visual question answering (VQA) formats are widely used for category-level privacy evaluation [shi2024assessment, mireshghallah2023can, staab2023beyond, tomekcce2025private, gu2025mllmguard, zhang2024benchmarking, sun2024trustllm]. CONFAIDE [mireshghallah2023can] employs psychological theories to evaluate model’s alignment with human judgment when evaluating the flow of privacy-related data. Some researches [staab2023beyond, tomekcce2025private] demonstrate that models can infer personal attributes from inputs, thereby infringing on personal privacy, and construct a benchmark to quantify the privacy reasoning capabilities of models. To further refine the scope of privacy protection, MLLMGuard [gu2025mllmguard] categorizes privacy types into personal privacy, trade secret, and state secret, but the limited dataset scale restricts the reliable of the evaluation. MultiTrust [zhang2024benchmarking] and TrustLLM [sun2024trustllm] assess privacy security of models from two dimensions: privacy awareness and privacy leakage, which improve the privacy assessment framework. As shown in Table I, we further expand the privacy scopes and dataset scale, constructing a more comprehensive evaluation benchmark. Furthermore, we restructure the evaluation framework with awareness and leakage by categorizing privacy leakage into granular subdimensions aligned with LVLM’s core competencies (perception, reasoning, and memorization). Finally, we design Insensitive Questions to evaluate the unintended impact of privacy protection mechanisms on benign query responsiveness.
III Task Definition
III-A Definition of Privacy
@@
Past category-level privacy benchmarks lack a formalized privacy definition. Some category-level privacy benchmarks [mireshghallah2023can, tomekcce2025private, zhang2024benchmarking] focus on the security of personal identifiable information, which is only a subset of privacy. MLLMGUARD [gu2025mllmguard] expands the scope of privacy to personal privacy, trade and state secrets, but it still lacks a clear definition of privacy.
@@
We define privacy formally as proprietary information confined to specific domains, characterized by: (a) general unknowability to extra-domain entities, (b) the condition that its acquisition by external parties constitutes ethical or legal violation, as shown below:
@@
where $G$ represents all groups, $P_{g}$ is the set of sensitive data for group $g$, $K(x)$ is the knowledge of group $x$ and $legal(g,k)$ represents that it is legal or ethical for group $g$ to get information $k$.
III-B Evaluation Objective
Our benchmark assesses the privacy risks of existing pre-trained LVLMs during inference. In our evaluation, rejection of sensitive queries constitutes effective privacy protection, while compliance with such requests is classified as a privacy breach.
Considering that a single evaluation benchmark may be inadequate to assess the privacy risks of LVLMs across various user groups, we focus on general users without special status, aiming to prevent them from extracting domain-specific sensitive information via LVLMs. For general users without special status, we define their privacy scope as the union of all group privacy settings, represented as:
@@ III-B Evaluation Objective
IV Multi-PA Benchmark
Figure 3: VQA Generation Process. For security reasons, we obscure the private parts in images. We build Image Database and Attribute Database by collecting images and designing attributes for each privacy category. For each task, we create a variety of question templates which will be randomly selected to generate samples. Each VQA sample is the combination of an image from Image Database and a question from Question Templates. For Privacy Question Detection and Privacy InfoFlow Assessment, context of each sample is respectively from corresponding question in Privacy Leakage and sample in CONFAIDE [mireshghallah2023can].
IV-A Dataset Overview
@@
Privacy Question Detection. Privacy Question Detection assesses the model’s capacity to discern the privacy risks of input queries. Notably, an image with privacy-related visual cues may be paired with a benign question, while a privacy-unrelated image may be associated with a sensitive question (e.g., querying for a politician’s home address given their publicly available photo is undeniably inappropriate). Therefore, the model’s assessment on the privacy sensitive of input queries may be context-dependent. To comprehensively evaluate a model’s awareness on the privacy sensitivity of input questions, we construct a dataset comprising image-question pairs, where images may contain private visual cues or not, and questions may be either privacy-related or privacy-unrelated. The model is tasked with determining the privacy relevance of each query.
Privacy InfoFlow Assessment. Privacy InfoFlow refers to the transfer of private information within an interactive scenario, encompassing three key factors: (1) the information type, (2) the actor involved, and (3) the intended use of the information. For example, in the scenario “health information is collected by a doctor in order to diagnosis”, “health information” represents the information type, “doctor” is the actor, and “diagnosis” is the use. This task evaluates how well a model’s understanding of privacy InfoFlow aligns with human.
@@
Perception Leakage. “Perception” Leakage assesses the risk of LVLMs extracting privacy present in input images. This leakage mode occurs in scenarios where attackers automate the extraction of private data. For example, attackers may gather large volumes of images from social platforms, a few of which may inadvertently contain private data (e.g., technology professionals might capture proprietary code in workplace photos shared online). Attackers could leverage the perceptual capabilities of LVLMs to automatically and cheaply extract such private data within images at scale. To evaluate the model’s vulnerability to leaking perceptible privacy, we design a set of perceivable privacy attributes for each privacy category. Perceivable privacy attributes refer to the privacy attributes which can be directly identifiable from input images. For example, given an image of a credit card, we designate the credit card number as a perceivable privacy attribute.
@@
Reasoning Leakage. “Reasoning” Leakage evaluates the risk of LVLMs inferring privacy based on visual cues within input images. Recent researches have demonstrated language models can infer privacy-sensitive data from indirect clues [staab2023beyond, tomekcce2025private]. Motivated by these findings, we consider the ability of models to infer privacy attributes as a crucial aspect of model’s privacy risks. To prevent models from facilitating the illicit prediction of private data, a privacy-preserving model should avoid generalizing its reasoning abilities to inferring privacy attributes. Therefore, for each privacy category, we curate a series of inferable privacy attributes for each privacy category. These attributes can be indirectly deduced from cues present in input images. For example, in an image containing a company’s revenue data, the company’s financial status might be considered as an inferable attribute.
@@
Memory Leakage. “Memory” Leakage evaluates the risk of LVLMs revealing privacy retained from its training process. Researches have shown that models can memorize training data and unintentionally disclose it during interactions [carlini2022quantifying, carlini2021extracting, jayaraman2022active, yu2023bag, staab2023beyond]. Privacy with long-term sensitivity, like PII of public figures, may be unintentionally exposed on social platforms and subsequently incorporated into model training datasets. The privacy leakage of such information may lead to detrimental consequences. To assess the risk of a model leaking private information from its training dataset, we define weakly associated privacy attributes for each privacy category. Weakly associated privacy attributes refer to information that cannot be directly or indirectly derived from the input data. However, models may still output these attributes based on private information it has memorized from its training dataset. For example, given images of some businessmen, we ask models about their home addresses. Such private data may have appeared in journalists’ reports. When these reports are used for model training, models might memorize such private data and provide correct answer during inference.
IV-B Dataset Construction